# Atomizer: Generalizing to new modalities by breaking satellite images down to a set of scalars

### v1

@@ Abstract
[…]
@@ 1 Introduction
Earth observation satellites capture our planet through different lenses, each offering distinct spectral, spatial, and temporal configurations. Many satellites are being launched with diverse spectral, spatial and temporal resolutions. While the first Earth observation satellite has been launched in 1972, more than half of them have been launched since 2020 (with 105 planned launches in 2025)usgscomp. Some satellites prioritize rich spectral information across dozens of bands but sacrifice spatial detail (e.g., MODIS), while others offer fine-grained spatial resolution with fewer spectral bands (e.g., Sentinel-2). These particularities are double-edged: they provide complementary insights for applications such as drought monitoring li2025improved, where MODIS’s frequent coverage is combined with Sentinel-2’s spatial detail to detect field-scale crop stress, as well as plant species classification pazos2024planted and crop mapping skakun2017combined. Yet traditional vision models, designed for uniform input formats, struggle to accommodate this heterogeneity. Recent work has begun to incorporate prior knowledge such as resolution or wavelengths into learning pipelines astruc2024anysat; reed2023scale. These methods often embed spatial resolution and spectral information through specialized modules within their model architectures. While these methods are suitable for some aspects of heterogeneity, most of them still rely on modality-specific encoders and patch representations, limiting their flexibility to unseen sensors configurations, often requiring costly pre-training steps li2025fleximo.
@@
We propose a different solution: instead of proposing a specific module for each type of input modality, we adapt the representation itself. Atomizer introduces a unified token-based framework that decomposes each observation into its most elementary unit, the reflectance of the band of a given pixel. Each token is enriched with acquisition-specific metadata, including spatial resolution, central wavelength, and spectral bandwidth. This atomic representation preserves all contextual information while enabling a simple, consistent framework which can be applied to a new optical modality that has a different combination of resolution, image size and spectral bands than those seen at training time. Atomizer builds on Perceiver jaegle2021perceiver, taking the core idea, the breaking down of images to a set of pixels, to the extreme, by mapping the set of scalars that compose an image into a compact latent space via cross-attention. This enables the model to handle inputs of arbitrary spatial size and channel depth. Metadata is encoded using structured strategies: Fourier features tancik2020fourier for positional encodings, and radial basis functions for spectral properties. This flexible framework allows each type of prior knowledge to be explicitly represented, making it straightforward to incorporate expert domain insights into the encoding process. As a result, Atomizer can reason across heterogeneous satellite observations without requiring architectural changes or retraining for new modalities.
@@ 2 Related Work
Addressing Resolution Variability. A limitation of most computer vision approaches for Earth observation, such as those based on convolutional neural networks (CNN) or Vision Transformers (ViT) dosovitskiy2020image, is their inability to effectively handle varying Ground Sample Distances (GSDs) at inference time. The fixed patch size selected during pre-training forces a trade-off. In high-resolution images, small patches lead to significant computational overhead, while in low-resolution, larger patches risk discarding critical fine-grained details. Several recent methods have attempted to address this resolution variability by incorporating GSD as prior knowledge. SenPa-MAE prexl2024senpa incorporates resolution as an explicit feature vector generated through multi-layer perceptrons, which is then added to the patch embeddings. Other approaches encode resolution information directly within the positional embeddings associated with each patch; this approach is taken by ScaleMAE reed2023scale; tseng2025galileo; astruc2024anysat, allowing these models to adapt feature extraction based on input resolution. FlexiMo li2025fleximo takes a different approach by applying bilinear interpolation to both raw image patches and patch embedding weights, enabling dynamic adjustment to varying spatial resolutions. While these methods improve performance across different GSD configurations, they remain constrained by their reliance on rigid patch-based tokenization.
@@
Incorporating Spectral Heterogeneity. Unlike natural images, satellite data contains spectral information that varies significantly across sensors liu2015survey, with different bands capturing distinct portions of the electromagnetic spectrum defined by their central wavelengths and bandwidths. Therefore, incorporating spectral configuration as prior knowledge is essential for effectively processing heterogeneous remote sensing data. Several approaches have emerged to address this challenge. FlexiMo li2025fleximo incorporates spectral information by using central wavelength to generate convolution kernels dynamically, enabling adaptation to varying channel configurations. Galileo tseng2025galileo splits input bands into semantically cohesive groups (e.g., RGB bands in Sentinel-2), enabling the network to model relationships between spectrally related channels more effectively. However, this approach introduces additional design complexity in determining optimal channel groupings and lacks flexibility when confronted with novel spectral configurations not encountered during training. SenPa-MAE prexl2024senpa embeds spectral responses as feature vectors added to patch embeddings, akin to positional encodings. While these methods improve handling of spectral variability, they typically require complex architectural modifications and still struggle to process truly heterogeneous observations with arbitrary band combinations without retraining.
@@
Handling Variable Spatial Dimensions. Beyond resolution and spectral variations, remote sensing models must also handle input images of varying spatial dimensions. Traditional vision transformers require fixed-size inputs, necessitating resizing or cropping that can distort spatial relationships or lose information. Several approaches have been proposed to address this limitation. FlexiMo and Galileo li2025fleximo; tseng2025galileo simply resize the patch embedding weights to fit with the dimensions of the input. AnySat astruc2024anysat takes a different approach by using fixed patch sizes that represent consistent ground distances in meters rather than pixel counts, allowing the model to maintain physical scale awareness across different resolution images. As a more significant departure from patch-based paradigms, Presto tseng2023lightweight eliminates patches entirely by encoding each pixel individually. While these approaches show strong results for handling varying input shapes, they typically require complex architectural modifications or introduce computational inefficiencies when processing high-resolution inputs.
@@
Towards Modality-Agnostic Architectures. While the approaches described above make progress in handling specific aspects of satellite data, most of them rely on complex modules added within the ViT framework, gradually increasing architectural complexity with each new modality characteristic. This approach contradicts the principle of Occam’s razor that simpler solutions are preferable. Instead of further complicating existing architectures, we draw inspiration from Perceiver jaegle2021perceiver, a model explicitly designed to process any modality: audio, video, images, through a unified framework. Perceiver addresses the computational scaling challenges of attention mechanisms by mapping arbitrary-sized inputs to a compact latent representation through cross-attention, enabling it to handle diverse input types without specialized encoders. By enriching inputs with modality-specific metadata through Fourier features tancik2020fourier rather than architectural modifications, Perceiver maintains a consistent architecture across modalities. Our approach leverages this insight while introducing a tokenization scheme specifically designed for the unique challenges of remote sensing data.
@@ 3 Methodology
Rather than developing increasingly complex architectural adaptations to accommodate heterogeneous satellite data, we propose a different approach that rethinks the input representation itself. Inspired by the Perceiver model jaegle2021perceiver, we bypass the need for fixed input formats by representing each observation as a fine-grained token. Specifically, we construct a token for each band of each pixel. These tokens include not only the raw band value, but also metadata such as spatial resolution, wavelength, and bandwidth, capturing both the content and context of each observation. To encode these attributes, we employ two different basis functions: (i) Fourier features for the band of value (e.g. reflectance) and the positional encoding; (ii) Radial basis functions for spectral properties such as wavelength and bandwidth. Our token-based design decouples the model from constraints such as fixed input shapes or specific acquisition schedules. By feeding the model a flat list of tokens, we can process input images of arbitrary resolution, spatial extent and spectral composition. This enables training a single encoder across diverse remote sensing datasets without requiring resampling, interpolation, or modality-specific adaptations.
@@ 3.1 Token Construction
Fourier Features. To encode continuous scalar metadata (such as time and resolution), we use Fourier features, also known as sinusoidal position encodings tancik2020fourier. These project a scalar input into a higher-dimensional space using sine and cosine functions, allowing the model to capture both low- and high-frequency variations.
Let $\tilde{x,}\in[0,1]$ be a normalized scalar,$L$ the number of frequency components and $f_{\text{max}}$ the maximum frequency component. The Fourier embedding function $\gamma(\cdot)$ is defined as:
$\gamma(\tilde{x};L,f_{\text{max}})=\left[\sin(\pi f_{1}\tilde{x}),\cos(\pi f_{1}\tilde{x}),\dots,\sin(\pi f_{L}\tilde{x}),\cos(\pi f_{L}\tilde{x})\right]$
@@
Spatial resolution encoding $\phi_{\text{res}}$: To encode spatial resolution and position information, we implement a resolution-aware positional encoding scheme. This approach builds upon Fourier features while incorporating ground sampling distance (GSD) as an explicit factor in the encoding. For each pixel at coordinates $(x,y)$ with resolution $g$, we first normalize the coordinates relative to the image center:
$x_{d}=\frac{x-x_{\text{center}}}{w/2},\quad y_{d}=\frac{y-y_{\text{center}}}{h/2}$
Where $w$ and $h$ are the width and height of the image, resulting in normalized coordinates $x_{d},y_{d}\in[-1,1]$. We then compute the resolution-modulated Fourier features for each coordinate as $\phi_{\text{res}}(x_{d},y_{d},g)=\text{Concat}(\gamma(x_{d}\frac{g}{G};L,f_{\text{max}}),\gamma(y_{d}\frac{g}{G};L,f_{\text{max}}))$, where $G$ is a reference GSD value chosen as a normalization constant, $g$ is the ground sampling distance (resolution) of the current band in meters per pixel, and $f_{k}$ is a set of frequencies linearly spaced between 1 and $f_{\text{max}}$. The ratio $\frac{g}{G}$ modulates the frequency based on resolution, allowing the model to differentiate between fine and coarse resolution inputs. We concatenate the original position of the pixel to the resulting positional encoding. This encoding yields a vector of size $12K+$ 1, where $K$ is the number of frequency components. The modulation of frequencies by the resolution ratio $\frac{g}{G}$ enables the model to represent the same spatial location differently based on resolution, providing an implicit understanding of scale. At finer resolutions (small $g$), the encoding varies more gradually across the image, while at coarser resolutions (large $g$), the encoding changes more rapidly, reflecting the different levels of detail available at each resolution.
@@
Spectral encoding $\phi_{\lambda}$: To encode the spectral characteristics of each band, defined by its central wavelength $\lambda$ and bandwidth $\mu$. We use a radial basis function (RBF) encoding scheme based on Gaussian kernels.
We define a set of $k$ Gaussian functions over the spectral domain:
$G=\left\{\mathcal{N}(\mu_{i},\sigma_{i})\;;\;i=0,\dots,k\right\}$
@@
Unlike typical uniform RBF placements, our Gaussians are strategically distributed: we allocate more narrow basis functions in spectral regions where many sensors and modalities operate (e.g., between 400 nm and 800 nm in the visible range). Fewer and wider Gaussians are used in sparsely populated spectral zones. This non-uniform distribution ensures higher resolution where it’s most needed, enabling the model to distinguish between closely spaced bands. Given a band defined by a central wavelength $\lambda$ and bandwidth $\mu$, we compute its spectral support as:
$P_{\text{min}}=\lambda-\frac{\mu}{2},\quad P_{\text{max}}=\lambda+\frac{\mu}{2}$
We uniformly sample wavelengths within this interval to form the set $P$. For each Gaussian $i$, we compute its maximum activation over $P$:
$\text{features}_{i}=\max_{\lambda^{\prime}\in P}\exp\left(-\frac{(\lambda^{\prime}-\mu_{i})^{2}}{2\sigma_{i}^{2}}\right)$
This yields a $k$-dimensional feature vector where each component reflects how much the band overlaps with a particular Gaussian. The final spectral encoding is then normalized:
$\phi_{\lambda}(\lambda,\mu)=\frac{\text{features}(\lambda,\mu)}{\|\text{features}(\lambda,\mu)\|_{2}}$
@@ 3.2 Architecture
The encoded tokens are processed using a Perceiver-style architecture jaegle2021perceiver. Our model ingests the unordered set of tokens and maps them into a compact latent representation through cross-attention mechanisms. Specifically, a set of $L$ learnable latent tokens attends to the input tokens, capturing information from the entire token set regardless of its size or structure.
To generate prediction logits from this latent space, we employ attention pooling as described in chen2023self, where the latent vectors are aggregated via attention mechanisms to produce classification outputs.
@@ 4 Experimental Framework
To evaluate Atomizer’s ability to generalize to unseen sensor configurations, we design a modality-disjoint evaluation protocol using the BigEarthNet dataset clasen2024reben, a large-scale benchmark for remote sensing image classification containing Sentinel-2 multispectral imagery across 19 land cover classes. Our experimental setup comprises 69,373 images for training, 65,618 for validation, and 63,972 for testing, with the complete dataset and modality configurations to be made publicly available on GitHub. We define distinct modalities by systematically varying three attributes: (1) spatial dimensions (pixel size), (2) ground sampling distance (GSD), and (3) spectral band composition (subset of Sentinel-2 bands). Each modality represents a unique combination of these attributes, simulating different satellite sensor configurations. This approach allows us to evaluate how models perform when encountering new sensor characteristics not seen during training. Our experimental design ensures that modalities used for training and testing are completely disjoint. We define separate training and evaluation modalities, with each training image associated exclusively with one training modality. By ensuring no image appears under multiple modalities during training, we prevent the model from explicitly learning relationships between different modality configurations. Test images are associated with entirely unseen modalities, creating a true cross-modality generalization challenge. Figure 2 illustrates the distinct modalities used for training and testing, highlighting their varying characteristics in terms of resolution, spatial dimensions, and spectral composition.
@@
Atomizer was trained for 40 epochs using a learning rate schedule consisting of a 5-epoch linear warmup followed by cosine annealing to zero. We used a batch size of 1024 distributed across two NVIDIA H100 GPUs. The model architecture consists of 4 cross-attention blocks, each containing 4 self-attention layers. To improve parameter efficiency, we employed weight sharing across all cross-attention blocks except the first. For classification output, we implemented latent attention pooling following chen2023self. To reduce the computational overhead induced by the cross attention layer, we randomly apply pruning on 50% of the model input tokens before each cross attention block.

### v2

@@ Abstract
Hugo Riffaud de Turckheim]hugo.riffaud–de-turckheim@inria.fr
@@ 1 Introduction
Earth observation satellites capture our planet through different lenses, each offering distinct spectral, spatial, and temporal configurations. Many satellites are being launched with diverse spectral, spatial and temporal resolutions. While the first Earth observation satellite has been launched in 1972, more than half of them have been launched since 2020 (with 105 planned launches in 2025)usgscomp (1). Some satellites prioritize rich spectral information across dozens of bands but sacrifice spatial detail (e.g., MODIS), while others offer fine-grained spatial resolution with fewer spectral bands (e.g., Sentinel-2). These particularities are double-edged: they provide complementary insights for applications such as drought monitoring li2025improved (2), where MODIS’s frequent coverage is combined with Sentinel-2’s spatial detail to detect field-scale crop stress, as well as plant species classification pazos2024planted (3) and crop mapping skakun2017combined (4). Yet traditional vision models, designed for uniform input formats, struggle to accommodate this heterogeneity. Recent work has begun to incorporate prior knowledge such as resolution or wavelengths into learning pipelines astruc2024anysat (5, 6). These methods often embed spatial resolution and spectral information through specialized modules within their model architectures. While these methods are suitable for some aspects of heterogeneity, most of them still rely on modality-specific encoders and patch representations, limiting their flexibility to unseen sensors configurations, often requiring costly pre-training steps li2025fleximo (7).
@@
We propose a different solution: instead of proposing a specific module for each type of input modality, we adapt the representation itself. Atomizer introduces a unified token-based framework that decomposes each observation into its most elementary unit, the reflectance of the band of a given pixel. Each token is enriched with acquisition-specific metadata, including spatial resolution, central wavelength, and spectral bandwidth. This atomic representation preserves all contextual information while enabling a simple, consistent framework which can be applied to a new optical modality that has a different combination of resolution, image size and spectral bands than those seen at training time. Atomizer builds on Perceiver jaegle2021perceiver (8), taking the core idea, the breaking down of images to a set of pixels, to the extreme, by mapping the set of scalars that compose an image into a compact latent space via cross-attention. This enables the model to handle inputs of arbitrary spatial size and channel depth. Metadata is encoded using structured strategies: Fourier features tancik2020fourier (9) for positional encodings, and radial basis functions for spectral properties. This flexible framework allows each type of prior knowledge to be explicitly represented, making it straightforward to incorporate expert domain insights into the encoding process. As a result, Atomizer can reason across heterogeneous satellite observations without requiring architectural changes or retraining for new modalities.
@@ 2 Related Work
Addressing Resolution Variability. A limitation of most computer vision approaches for Earth observation, such as those based on convolutional neural networks (CNN) or Vision Transformers (ViT) dosovitskiy2020image (10), is their inability to effectively handle varying Ground Sample Distances (GSDs) at inference time. The fixed patch size selected during pre-training forces a trade-off. In high-resolution images, small patches lead to significant computational overhead, while in low-resolution, larger patches risk discarding critical fine-grained details. Several recent methods have attempted to address this resolution variability by incorporating GSD as prior knowledge. SenPa-MAE prexl2024senpa (11) incorporates resolution as an explicit feature vector generated through multi-layer perceptrons, which is then added to the patch embeddings. Other approaches encode resolution information directly within the positional embeddings associated with each patch; this approach is taken by ScaleMAE reed2023scale (6, 12, 5), allowing these models to adapt feature extraction based on input resolution. FlexiMo li2025fleximo (7) takes a different approach by applying bilinear interpolation to both raw image patches and patch embedding weights, enabling dynamic adjustment to varying spatial resolutions. While these methods improve performance across different GSD configurations, they remain constrained by their reliance on rigid patch-based tokenization.
@@
Incorporating Spectral Heterogeneity. Unlike natural images, satellite data contains spectral information that varies significantly across sensors liu2015survey (13), with different bands capturing distinct portions of the electromagnetic spectrum defined by their central wavelengths and bandwidths. Therefore, incorporating spectral configuration as prior knowledge is essential for effectively processing heterogeneous remote sensing data. Several approaches have emerged to address this challenge. FlexiMo li2025fleximo (7) incorporates spectral information by using central wavelength to generate convolution kernels dynamically, enabling adaptation to varying channel configurations. Galileo tseng2025galileo (12) splits input bands into semantically cohesive groups (e.g., RGB bands in Sentinel-2), enabling the network to model relationships between spectrally related channels more effectively. However, this approach introduces additional design complexity in determining optimal channel groupings and lacks flexibility when confronted with novel spectral configurations not encountered during training. SenPa-MAE prexl2024senpa (11) embeds spectral responses as feature vectors added to patch embeddings, akin to positional encodings. While these methods improve handling of spectral variability, they typically require complex architectural modifications and still struggle to process truly heterogeneous observations with arbitrary band combinations without retraining.
@@
Handling Variable Spatial Dimensions. Beyond resolution and spectral variations, remote sensing models must also handle input images of varying spatial dimensions. Traditional vision transformers require fixed-size inputs, necessitating resizing or cropping that can distort spatial relationships or lose information. Several approaches have been proposed to address this limitation. FlexiMo and Galileo li2025fleximo (7, 12) simply resize the patch embedding weights to fit with the dimensions of the input. AnySat astruc2024anysat (5) takes a different approach by using fixed patch sizes that represent consistent ground distances in meters rather than pixel counts, allowing the model to maintain physical scale awareness across different resolution images. As a more significant departure from patch-based paradigms, Presto tseng2023lightweight (14) eliminates patches entirely by encoding each pixel individually. While these approaches show strong results for handling varying input shapes, they typically require complex architectural modifications or introduce computational inefficiencies when processing high-resolution inputs.
@@
Towards Modality-Agnostic Architectures. While the approaches described above make progress in handling specific aspects of satellite data, most of them rely on complex modules added within the ViT framework, gradually increasing architectural complexity with each new modality characteristic. This approach contradicts the principle of Occam’s razor that simpler solutions are preferable. Instead of further complicating existing architectures, we draw inspiration from Perceiver jaegle2021perceiver (8), a model explicitly designed to process any modality: audio, video, images, through a unified framework. Perceiver addresses the computational scaling challenges of attention mechanisms by mapping arbitrary-sized inputs to a compact latent representation through cross-attention, enabling it to handle diverse input types without specialized encoders. By enriching inputs with modality-specific metadata through Fourier features tancik2020fourier (9) rather than architectural modifications, Perceiver maintains a consistent architecture across modalities. Our approach leverages this insight while introducing a tokenization scheme specifically designed for the unique challenges of remote sensing data.
@@ 3 Methodology
Rather than developing increasingly complex architectural adaptations to accommodate heterogeneous satellite data, we propose a different approach that rethinks the input representation itself. Inspired by the Perceiver model jaegle2021perceiver (8), we bypass the need for fixed input formats by representing each observation as a fine-grained token. Specifically, we construct a token for each band of each pixel. These tokens include not only the raw band value, but also metadata such as spatial resolution, wavelength, and bandwidth, capturing both the content and context of each observation. To encode these attributes, we employ two different basis functions: (i) Fourier features for the band of value (e.g. reflectance) and the positional encoding; (ii) Radial basis functions for spectral properties such as wavelength and bandwidth. Our token-based design decouples the model from constraints such as fixed input shapes or specific acquisition schedules. By feeding the model a flat list of tokens, we can process input images of arbitrary resolution, spatial extent and spectral composition. This enables training a single encoder across diverse remote sensing datasets without requiring resampling, interpolation, or modality-specific adaptations.
@@ 3.1 Token Construction
Fourier Features. To encode continuous scalar metadata (such as time and resolution), we use Fourier features, also known as sinusoidal position encodings tancik2020fourier (9). These project a scalar input into a higher-dimensional space using sine and cosine functions, allowing the model to capture both low- and high-frequency variations.
Let $\tilde{x,}\in[-1,1]$ be a normalized scalar,$L$ the number of frequency components and $f_{\text{max}}$ the maximum frequency component. The Fourier embedding function $\gamma(\cdot)$ is defined as:
[…]
@@
Spatial resolution encoding $\phi_{\text{res}}$: To encode spatial resolution and position information, we implement a resolution-aware positional encoding scheme that incorporates ground sampling distance (GSD) as an explicit factor in Fourier features.
For each pixel at coordinates $(x,y)$ with resolution $g$, we compute resolution-scaled coordinates that span the physical extent of the image:
$x_{\text{scaled}}=\left(x-\frac{w}{2}\right)\cdot\frac{g}{G},\quad y_{\text{scaled}}=\left(y-\frac{h}{2}\right)\cdot\frac{g}{G}$
where $w$ and $h$ are the width and height of the image in pixels, $g$ is the ground sampling distance in meters per pixel, and $G$ is a reference normalization constant. The resolution-modulated Fourier features are then computed as:
$\phi_{\text{res}}(x,y,g)=\text{Concat}\left(\gamma(x_{\text{scaled}};L,f_{\text{max}}),\gamma(y_{\text{scaled}};L,f_{\text{max}})\right)$
where $\gamma(\cdot)$ is the Fourier encoding function, $L$ is the number of frequency components, and $f_{\text{max}}$ is the maximum frequency.
The scaling factor $\frac{g}{G}$ directly modulates the coordinate space based on physical resolution. At finer resolutions (small $g$), coordinates span a smaller range corresponding to a smaller physical area, while at coarser resolutions (large $g$), coordinates span a larger range reflecting greater physical coverage. This enables the model to maintain consistent spatial understanding across different sensor configurations. The encoding yields a vector of size $4L+2$, where we concatenate the original pixel coordinates to preserve absolute positional information.
@@
Spectral encoding $\phi_{\lambda}$: To encode the spectral characteristics of each band, defined by its central wavelength $\lambda$ and bandwidth $\mu$, we use a radial basis function (RBF) encoding scheme based on Gaussian kernels. We use radial basis functions $\phi_{\lambda}(\lambda,\mu)$ to encode spectral characteristics. Each band’s spectral support (defined by central wavelength $\lambda$ and bandwidth $\mu$) is matched against strategically placed Gaussians.
$\phi_{\lambda}(\lambda,\mu)=\left[\int_{\lambda-\mu/2}^{\lambda+\mu/2}\mathcal{N}(\mu_{1},\sigma_{1})d\lambda^{\prime},\ldots,\int_{\lambda-\mu/2}^{\lambda+\mu/2}\mathcal{N}(\mu_{k},\sigma_{k})d\lambda^{\prime}\right]$
@@
Unlike typical uniform RBF placements, our Gaussians are strategically distributed: we allocate more narrow basis functions in spectral regions where many sensors and modalities operate (e.g., between 400 nm and 800 nm in the visible range). Fewer and wider Gaussians are used in sparsely populated spectral zones. This non-uniform distribution ensures higher resolution where it’s most needed, enabling the model to distinguish between closely spaced bands. For each Gaussian $i$, we compute the integral of the Gaussian function over the band’s spectral support range $\lambda-\mu/2,\lambda+\mu/2]$. This yields a $k$-dimensional feature vector where each component reflects how much the band overlaps with a particular Gaussian. The final spectral encoding is then L2-normalized.
@@ 3.2 Architecture
The encoded tokens are processed using a Perceiver-style architecture jaegle2021perceiver (8). Our model ingests the unordered set of tokens and maps them into a compact latent representation through cross-attention mechanisms. Specifically, a set of $L$ learnable latent tokens attends to the input tokens, capturing information from the entire token set regardless of its size or structure.
To generate prediction logits from this latent space, we employ attention pooling as described in chen2023self (15), where the latent vectors are aggregated via attention mechanisms to produce classification outputs.
@@ 4 Experimental Framework
To evaluate Atomizer’s ability to generalize to unseen sensor configurations, we design a modality-disjoint evaluation protocol using the BigEarthNet dataset clasen2024reben (16), a large-scale benchmark for remote sensing image classification containing Sentinel-2 multispectral imagery across 19 land cover classes. Our experimental setup comprises 69,373 images for training, 65,618 for validation, and 63,972 for testing, with the complete dataset and modality configurations to be made publicly available on GitHub. We define distinct modalities by systematically varying three attributes: (1) spatial dimensions (pixel size), (2) ground sampling distance (GSD), and (3) spectral band composition (subset of Sentinel-2 bands). Each modality represents a unique combination of these attributes, simulating different satellite sensor configurations. This approach allows us to evaluate how models perform when encountering new sensor characteristics not seen during training. Our experimental design ensures that modalities used for training and testing are completely disjoint. We define separate training and evaluation modalities, with each training image associated exclusively with one training modality. By ensuring no image appears under multiple modalities during training, we prevent the model from explicitly learning relationships between different modality configurations. Test images are associated with entirely unseen modalities, creating a true cross-modality generalization challenge. Figure 2 illustrates the distinct modalities used for training and testing, highlighting their varying characteristics in terms of resolution, spatial dimensions, and spectral composition.
@@
Atomizer was trained for 40 epochs using a learning rate schedule consisting of a 5-epoch linear warmup followed by cosine annealing to zero. We used a batch size of 1024 distributed across two NVIDIA H100 GPUs. The model architecture consists of 4 cross-attention blocks, each containing 4 self-attention layers. To improve parameter efficiency, we employed weight sharing across all cross-attention blocks except the first. For classification output, we implemented latent attention pooling following chen2023self (15). To reduce the computational overhead induced by the cross attention layer, we randomly apply pruning on 50% of the model input tokens before each cross attention block.